A Bibliography of Books about the Internet and Networking: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 May 2024 Version 1.03

Title word cross-reference

Administration [FM20].

Course [MFD20]. Creation [FM20].

First [MFD20]. Freedom [Fed24].

Identifiers [FM20]. Internet [Fed24].

Network [MFD20].

Open [Fed24].

Restoring [Fed24].

Safeguarding [Fed24]. Science [MFD20].

Securing [Fed24].

Unique [FM20].

References

Commission:2024:SSO

[Fed24] Federal Communications Commission. Safeguarding and securing the Open Internet and restoring Internet freedom. Ruling FCC 24-52, Federal Communications Commission, Washington, DC 20554, USA, May 7, 2024. 513 (final 82+ missing: should

be 595+) pp. URL https://docs.
fcc.gov/public/attachments/FCC-
24-52A1.pdf.

Fidler:2020:CAU

[FM20] Bradley Fidler and Russ Mundy.

The creation and administration of unique identifiers, 1967–2017. Internet Corporation for Assigned Names and Numbers (ICANN) report, ????, November 2020. 118

pp. URL https://www.icann.org/
en/system/files/files/creation-
administration-unique-identifiers-
1967-2017-18nov20-en.pdf.

Menczer:2020:FCN

[MFD20] Filippo Menczer, Santo Fortunato, and Clayton A. Davis. A First Course in Network Science. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-47113-7 (hardcover), 1-108-65394-4 (e-book). xv + 258 pp. LCCN TK5105.5 .M46 2020, T57.85 .M46 2020.